the average of 5 test scores is a 78. what is the fifth score if the first 4 scores are 71, 92, 68, and 84?
Answer by w_parminder(53)

Average of 5 test scores = 78
So sum of 5 test scores = 78*5 = 390
Sum of first 4 test scores = 71+92+68+84 = 315
So the fifth score = 390-315 = 75